FREE TROPOSPHERIC MEASUREMENTS OF BIOMASS BURNING EMISSIONS BY THE ATMO-SPHERIC CHEMISTRY EXPERIMENT (ACE)

\underline{CURTIS P. RINSLAND}, NASA Langley Research Center Mail Stop 401A, Hampton, VA 23681 U.S.A. (c.p.rinsland@larc.nasa.gov); CHRIS D. BOONE, B. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada (cboone@acdbox.uwaterloo.ca and bernath@uwaterloo.ca); LINDA CHIOU, Systems and Applied Sciences Corporation, Hampton, VA U.S.A (l.s.chiou@larc.nasa.gov).

The Atmospheric Chemistry Experiment (ACE) was successfully launched in August 2003 to monitor the chemical composition Earth’s atmosphere with measurements obtained primarily in solar occultation mode. The primary instrument is a Fourier transform spectrometer (FTS) covering the infrared from 750-4400 cm^{−1} simultaneously with a maximum optical path difference of 25 cm (0.02 cm^{−1} resolution). We describe free tropospheric measurements of biomass burning products such as HCN, CO, C_{2}H_{6}, CH_{3}Cl, HCOOH and 1.02 μm extinction and their interpretation based on other satellite measurements and back trajectory calculations.